OpenStudy (shamil98):

Solve the initial value problem: \[y'' + 3y = 0\] \[y(0) = 1 , y'(0)= 3\] \[r^2 + 3 = 0 \] \[r = \pm i √3\] Using Euler's equation: \[\alpha = -b/(2a) , \beta = \sqrt{4ac-b^2}\] \[\large e^{\alpha x}(c_1\cos \beta x+c_2\sin \beta x)\] \[\alpha = 0 , \beta = \sqrt{3}\] The general solution is: \[y(x) = c_1\cos \sqrt{3}~x +c_2\sin \sqrt{3}~x\] Am i right so far?..

OpenStudy (shamil98):

\[y'(x) = -c_1 \sin \sqrt{3}~x + c_2 \cos \sqrt{3}~x\] \[ 1 = c_1\] \[3 = c_2\] So: \[y(x) = \cos \sqrt{3}~x +3\sin \sqrt{3}~x\]

OpenStudy (math&ing001):

You forgot the sqrt(3) on the derivative, but other than that I don't see any mistakes.

OpenStudy (shamil98):

Where on the derivative?..

OpenStudy (math&ing001):

y'(x)=-c1*sqrt(3)*sin (sqrt(3)x) + c2*sqrt(3)*cos(sqrt(3)x)

OpenStudy (shamil98):

I could rewrite it like this if that's what you meant. \[y(x) = c_1\cos x\sqrt{3} +c_2\sin x\sqrt{3}\] \[y'(x) = -c_1 \sin x\sqrt{3}+ c_2 \cos x\sqrt{3}\] oh product rule, right, forgot about that .